OpenStudy (anonymous):

lim Δx→0+ ((5/x+Δx)-(5/x))/Δx

OpenStudy (anonymous):

Here, f(x)=5/x and you need to find its derivative that is f'(x)

OpenStudy (anonymous):

thing is, we haven't discussed derivatives yet

OpenStudy (anonymous):

geth the common denominator in the top part and then see what you get.

OpenStudy (anonymous):

(5Δx/x²+xΔx)/Δx

OpenStudy (anonymous):

u missed the negative sign

OpenStudy (anonymous):

now rewrite the complex fraction as multiplication, i.e., (a/b)/(c/d) = a/b * d/c

OpenStudy (anonymous):

and u can cancel delta x also..

OpenStudy (anonymous):

(5Δx/x²+xΔx) * (1/Δx)

OpenStudy (anonymous):

5Δx/Δx(x²+xΔx)

OpenStudy (anonymous):

dont you think you are missing negative sign?

OpenStudy (anonymous):

i cant seem to find where it is missing from

OpenStudy (anonymous):